OpenStudy (anonymous):

On the graph of the equation 6x – y = –1, what is the value of the y-intercept?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

is the answer 1?

OpenStudy (saifoo.khan):

y = 6x + 1

OpenStudy (saifoo.khan):

So, 1.

OpenStudy (anonymous):

y is 1

OpenStudy (katrinakaif):

Convert this into a y= mx + b format When you do so, 6x - y = -1 -y = -6x - 1 y = 6x + 1 The y - intercept is always 'b' resulting the y-intercept to be 1

OpenStudy (anonymous):

thank u

OpenStudy (katrinakaif):

Anytime =)